OpenStudy (anonymous):

list of rational numbers up to 50?

OpenStudy (anonymous):

not possible as there are an infinite number of them

OpenStudy (dan815):

infinite

OpenStudy (anonymous):

are you sure that is what the question asks?

OpenStudy (anonymous):

Oh sorry I meant up to the square root of 50. :) Also it's not really for a school question. I'm studying for high school.

OpenStudy (anonymous):

still not possible

OpenStudy (anonymous):

there are an infinite number of rational numbers between 0 and 1 or between any two numbers you cannot list them

OpenStudy (dan815):

satellite your bicycle for some reason is memorizing

OpenStudy (anonymous):

memorizing what?

OpenStudy (dan815):

mesmerizing

OpenStudy (anonymous):

oooh i bought that with my first 1,000 medals

OpenStudy (anonymous):

For example... √1 √2 √3 √4 √5 √6 √7 √8 √9 √10 √11 √12 √13 √14 √15 √16 √17 √18 √19 √20 Which ones are rational?

OpenStudy (anonymous):

oh i see the ones that are perfect square, and only those

OpenStudy (anonymous):

Perfect Square?

OpenStudy (anonymous):

for example \[\sqrt{16}=4\] is rational, but \(\sqrt{15}\) is not

OpenStudy (dan815):

i told u to read that proof sqrt prime = irrational

OpenStudy (anonymous):

\[\sqrt 1=1\] that is ratioal

OpenStudy (vishweshshrimali5):

But still there will be infinite rational numbers.

OpenStudy (anonymous):

\[\sqrt4=2\] is also rational

OpenStudy (anonymous):

@dan815 The link wouldn't work on my computer. @satellite73 So 19 would be irrational?

OpenStudy (anonymous):

think the question is really "which integers have rational square roots?"

OpenStudy (anonymous):

@CrazyCountryGirl yes they are all irrational, unless they are actually integers

OpenStudy (anonymous):

OK. :) Would 0.278 with a line above the 8 be rational?

OpenStudy (anonymous):

when i say "perfect square" i meant for example \(25\) because \(25=5^2\)

OpenStudy (anonymous):

yes, any repeating decimal is also rational

OpenStudy (anonymous):

and the line above the \(8\) means the decimal repeats

OpenStudy (dan815):

suppose Sqrt P = m/n <--- m and n being integers making it a rational number then P=m^2/n^2 and n^2*P=m^2 , but m^2 has an even number of prime factors and n^2*p = an odd number of prime factors therefore by contradiction sqrtp cannot equal m/n

OpenStudy (anonymous):

\[.027\overline{8}=.0278888888...\]

OpenStudy (anonymous):

any terminating decimal or repeating decimal is rational the decimal for an irrational number is non repeating, not terminating

OpenStudy (anonymous):

Oh.... So which of these would be rational, explain please. :) π 0.278 (line above the 8) –0.8764… √19 @satellite73

OpenStudy (anonymous):

all are irrational except the repeating decimal

OpenStudy (anonymous):

Oh so the third one?

OpenStudy (anonymous):

no, the second one is the repeating decimal that is the only rational number in that list

OpenStudy (anonymous):

did you figure out how to get youtube videos yet?

OpenStudy (anonymous):

Oh. :) Can you explain? Yes. I'm not allowed to use it on my mom s computer though. Last time I did, it got the blue screen of death. :(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol, you didn't have to do extra chores around the house. :(

OpenStudy (anonymous):

\(\pi\) is irrational because... well because it is that is why it is represented by a greek letter rather than a fraction

OpenStudy (anonymous):

\[\sqrt{19}\] is irrational because the square root of anything is irrational, unless the number inside is a perfect square

OpenStudy (anonymous):

you are supposed to think \[–0.8764…\] is irrational because of the \(...\) and no pattern, so it is not supposed to repeat (a stupid assumption, but that is what they want you to think)

OpenStudy (anonymous):

so the only rational one is the repeating decimal

OpenStudy (anonymous):

i wonder if Christina Grimmie has to do chores around the house when she makes youtube videos

OpenStudy (anonymous):

:P